# TerraMesh: A Planetary Mosaic of Multimodal Earth Observation Data

### v1

@@ Abstract
Large-scale foundation models in Earth Observation can learn versatile, label-efficient representations by leveraging massive amounts of unlabeled data. However, existing public datasets are often limited in scale, geographic coverage, or sensor variety. We introduce TerraMesh, a new globally diverse, multimodal dataset combining optical, synthetic aperture radar, elevation, and land-cover modalities in an Analysis-Ready Data format. TerraMesh includes over 9 million samples with eight spatiotemporal aligned modalities, enabling large-scale pre-training and fostering robust cross-modal correlation learning. We provide detailed data processing steps, comprehensive statistics, and empirical evidence demonstrating improved model performance when pre-trained on TerraMesh. The dataset will be made publicly available with a permissive license.
Figure 1: Samples from the TerraMesh dataset with seven spatiotemporal aligned modalities. Sentinel-2 L2A uses IRRG pseudo-coloring and Sentinel-1 RTC is visualized as VH-VV-VH. To account for their value ranges, VV is scaled from -30 – 5 db while VH is scaled from -40 – 0 db. Copernicus DEM is scaled based on the image value range with an additional 10 meter buffer to highlight flat scenes.
@@ 2 Related work
Table 2: Summary of modalities included in TerraMesh. The range includes the typical value range, including any inherent offsets. Every co-registered sample includes seven out of eight modalities due to the source datasets providing either S-1 GRD or S-1 RTC data.
@@
Sentinel-2 L1C Optical DN 1000 – 11000 13 Top-of-atmosphere reflectance
Sentinel-2 L2A Optical DN 1000 – 11000 12 Bottom-of-atmosphere reflectance
@@ 3.2 Data Processing
In assembling TerraMesh, we leveraged data from multiple repositories. MajorTOM-Core provides S-2 data from the CREODIAS platform and S-1 data from Planetary Computer [12], whereas SSL4EO-S12,v1.1 relies on Google Earth Engine for its S-1 and S-2 acquisitions [3]. We obtained the Copernicus DEM from the Copernicus Data Space EcosystemCopernicus DEM source: https://dataspace.copernicus.eu/explore-data/data-collections/copernicus-contributing-missions/collections-description/COP-DEM, and further acquired ESRI’s LULC maps from Planetary ComputerESRI LULC source: https://planetarycomputer.microsoft.com/dataset/io-lulc-annual-v02.
@@
A random 99%-1% train-validation split is applied to the selected MajorTOM-Core grid cells. Any SSL4EO-S12 v1.1 patch that intersects the validation extent is added to the validation set; further patches spatially overlapping the joined validation area are removed to mitigate data leakage. All Sentinel-2 bands are resampled to 10 m using nearest interpolation, and a +1000 offset is added to pre-2022 data, following ESA’s latest processing standard [10].
@@
We also provide NDVI layers derived from Sentinel-2 L2A, by removing any offsets, clipping negative values to zero, and adding a small epsilon in the denominator. Copernicus DEM data, acquired at 30 m resolution, is bilinearly resampled and projected to the Sentinel-2 grid. ESRI LULC products at 10 m resolution are downloaded for the closest available year from the S-2 timestamp and regridded correspondingly. Missing LULC pixels are assigned the no data class. To improve the alignment between the satellite images and the LULC maps, we replace the ESRI cloud class with no data and add the cloud and snow/ice predictions from the SEnSeI v2 masks.
@@
We finalize the dataset by shuffling and stacking patches into Zarr archives (ZipStore, version 2Zarr documentation: https://zarr-specs.readthedocs.io/en/latest/v2/v2.0.html) using xarray Datasets. Each file includes a chunk of 64 samples, similar to SSL4EO-S12 v1.1 [3]. This approach consolidates large numbers of samples into fewer individual files, which is easier to handle on HPC clusters and other file systems with a limited number of inodes. The Zarr files are fully compatible with xarray, support lazy loading through dask, and include aligned metadata. We optimize the data size by storing the Sentinel-2 bands and DEM as 16-bit integers, while Sentinel-1 and NDVI are saved as 16-bit floats; both LULC and Sentinel-2 RGB are stored in 8-bit format. In conjunction with Zarr’s built-in compression, these measures substantially reduce storage overhead without sacrificing data fidelity. When comparing a webdataset format using numpy and tar and the sample data types to Zarr, the dataset size reduces from 36.2 TB to only 14 TB, achieving a compression factor of 2.6. The compression factor varies between modalities, ranging from 20 for LULC to only 1.6 for S-1 and 1.4 for S-2 RGB. By total size, S-2 compression has the biggest effect, saving 17.9 TB with factors of 2.7 (S-2 L2A) and 3.0 (S-2 L1C).
@@ 3.3 Dataset description
TerraMesh includes 14 TB of compressed data and consists of 9 089 536 training and 89 088 validation samples, aggregated into 143 416 Zarr Zip files per modality for streamlined data access. Each sample features seven spatially aligned modalities, resulting in 64M individual patches, all modalities considered. Figure 3 illustrates the temporal distribution of these samples, using the S-2 timestamp. The data covers eight years with the majority of acquisitions between 2019 and 2023.
@@
Data loading performance becomes a central concern for smaller multimodal models, which often rely on rapid data retrieval. To address this, we store groups of 64 patches in a single Zarr chunk, enabling entire minibatches to be read in a single I/O operation. This strategy improves GPU utilization as 64 samples fit into the memory of most GPUs. Larger models, however, may require splitting these chunks to accommodate reduced effective batch sizes; the higher computational cost in such cases diminishes the impact of data loading for model training. Although fixing chunk sizes reduces the level of per-epoch shuffling at the minibatch level, global multi-GPU training typically provides sufficient stochastic diversity across the dataset.
@@ 6 Acknowledgement
We provide pseudocode for generating S-2 RGB images and NDVI maps from Sentinel-2 L2A in Listing LABEL:lst:rgb and LABEL:lst:ndvi. We refer to Subsection 3.2 for a detailed description.
On the following pages, we present additional examples from both data sources in Figure 6 and 7. All examples are taken from each subset’s first validation Zarr file, and we manually select representative samples to showcase the full range of landscapes. The focus on urban areas in SSL4EO-S12 is visible, which benefits the pre-training compared to homogeneous sampling. Because of the additional scaling for S2 RGB, the cloudy images are not very dark, while the clouds keep their structure and are not flattened to a single white value. Notice the alignment of the cloud and ice classes in the LULC data due to the augmentation using the SEnSeI v2 cloud and ice masks. We only annotate thick clouds as the land surface is still visible below thin cloud layers.
@@
import numpy as np def transform_rgb(rbg, offset=1000): rbg = rbg - offset Q2, Q98 = np.quantile(rbg, [0.02, 0.98]) rbg = np.where(rbg >= Q2, rbg, Q2 + (rbg - Q2) * 0.5) rbg = np.where(rbg <= Q98, rbg, Q98 + (rbg - Q98) * 0.5) Q02, Q50, Q998 = np.quantile(rbg, [0.002, 0.5, 0.998]) U = max(2000, Q998) L = 0 if Q50 < 1000 else Q02 rbg = (rbg - L) / (U - L) * 255 rbg = np.clip(rbg, 0, 255) return rbg.astype(np.uint8)
@@
def calculate_ndvi(b04, b08, offset=1000): b04 = (b04 - offset).clip(0) b08 = (b08 - offset).clip(0) ndvi = (b08 - b04) / (b08 + b04 + 1e-6) return ndvi
Figure 6: Additional examples from the MajorTOM-Core subset with seven spatiotemporal aligned modalities. Sentinel-2 L2A uses IRRG pseudo-coloring and Sentinel-1 RTC is visualized as VH-VV-VH. To account for their value ranges, VV is scaled from -30 – 5 db while VH is scaled from -40 – 0 db. Copernicus DEM is scaled based on the image value range with an additional 10 meter buffer.
Figure 7: Additional examples from the SSL4EO-S12 subset with seven spatiotemporal aligned modalities. Sentinel-2 L2A uses IRRG pseudo-coloring and Sentinel-1 RTC is visualized as VH-VV-VH. To account for their value ranges, VV is scaled from -30 – 5 db while VH is scaled from -40 – 0 db. Copernicus DEM is scaled based on the image value range with an additional 10 meter buffer.

### v2

@@ Abstract
Large-scale foundation models in Earth Observation can learn versatile, label-efficient representations by leveraging massive amounts of unlabeled data. However, existing public datasets are often limited in scale, geographic coverage, or sensor variety. We introduce TerraMesh, a new globally diverse, multimodal dataset combining optical, synthetic aperture radar, elevation, and land-cover modalities in an Analysis-Ready Data format. TerraMesh includes over 9 million samples with eight spatiotemporal aligned modalities, enabling large-scale pre-training. We provide detailed data processing steps, comprehensive statistics, and empirical evidence demonstrating improved model performance when pre-trained on TerraMesh. The dataset is hosted at https://huggingface.co/datasets/ibm-esa-geospatial/TerraMesh.
Figure 1: Samples from the TerraMesh dataset with seven spatiotemporal aligned modalities. Sentinel-2 L2A uses IRRG pseudo-coloring, and Sentinel-1 RTC is visualized in dB scale as VH-VV-VV/VH. Copernicus DEM is scaled based on the image value range with an additional 10 meter buffer to highlight flat scenes.
@@ 2 Related work
Table 2: Summary of modalities included in TerraMesh. The range includes the typical value range. Every co-registered sample includes seven out of eight modalities due to the source datasets providing either S-1 GRD or S-1 RTC data.
@@
Sentinel-2 L1C Optical DN 0 – 10000 13 Top-of-atmosphere reflectance
Sentinel-2 L2A Optical DN 0 – 10000 12 Bottom-of-atmosphere reflectance
@@ 3.2 Data Processing
In assembling TerraMesh, we leveraged data from multiple repositories. MajorTOM-Core provides S-2 data from the CREODIAS platform and S-1 data from Planetary Computer [12], whereas SSL4EO-S12 v1.1 relies on Google Earth Engine for its S-1 and S-2 acquisitions [3]. We obtained the Copernicus DEM from the Copernicus Data Space EcosystemCopernicus DEM source: https://dataspace.copernicus.eu/explore-data/data-collections/copernicus-contributing-missions/collections-description/COP-DEM, and further acquired ESRI’s LULC maps from Planetary ComputerESRI LULC source: https://planetarycomputer.microsoft.com/dataset/io-lulc-annual-v02.
@@
A random 99%-1% train-validation split is applied to the selected MajorTOM-Core grid cells. Any SSL4EO-S12 v1.1 patch that intersects the validation extent is added to the validation set; further patches spatially overlapping the joined validation area are removed to mitigate data leakage. All Sentinel-2 bands are resampled to 10 m using nearest interpolation, and the +1000 offset is removed from post-2022 data, which was added in ESA’s latest processing standard [10].
@@
We also provide NDVI layers derived from Sentinel-2 L2A, by clipping negative values to zero and adding a small epsilon in the denominator. Copernicus DEM data, acquired at 30 m resolution, is bilinearly resampled and projected to the Sentinel-2 grid. ESRI LULC products at 10 m resolution are downloaded for the closest available year from the S-2 timestamp and regridded correspondingly. Missing LULC pixels are assigned the no data class. To improve the alignment between the satellite images and the LULC maps, we replace the ESRI cloud class with no data and add the cloud and snow/ice predictions from the SEnSeI v2 masks.
@@
We finalize the dataset by saving each patch as a Zarr Zip file (ZipStore, version 2Zarr documentation: https://zarr-specs.readthedocs.io/en/latest/v2/v2.0.html) using xarray Datasets. The patches are shuffled and stored in tar archives that can be used with WebDatasetWebDataset documentation: https://github.com/webdataset/webdataset. This approach consolidates large numbers of samples into fewer individual files, which is easier to handle on HPC clusters and other file systems with a limited number of inodes. We further optimize the data size by storing the Sentinel-2 bands and DEM as 16-bit integers, while Sentinel-1 and NDVI are saved as 16-bit floats; both LULC and Sentinel-2 RGB are stored in 8-bit format. In conjunction with Zarr’s built-in compression, these measures substantially reduce storage overhead without sacrificing data fidelity. When comparing a TIF data format and Zarr, the dataset size reduces from 41 TB to only 16 TB, achieving a compression factor of 2.9. The compression factor varies between modalities, ranging from 20 for LULC to only 2.5 for S-2 RGB. By total size, S-2 compression has the biggest effect, saving 18.6 TB with factors of 2.6 (S-2 L2A) and 2.9 (S-2 L1C). Using WebDataset with Zarr Zip files also optimizes the data loading speed, as demonstrated in the supplementary material.
@@ 3.3 Dataset description
TerraMesh includes 16 TB of compressed data and consists of 9 089 536 training and 89 088 validation samples, aggregated into tar archives per modality for streamlined data access with WebDataset. Each spatiotemporal sample features seven aligned modalities, resulting in 64M individual patches, all modalities considered. Figure 3 illustrates the temporal distribution of these samples, using the S-2 timestamp. The data covers eight years with the majority of acquisitions between 2019 and 2023.
@@ 6 Acknowledgement
This supplementary material includes (1) pseudocode for generating the additonal Sentinel-2 modalities, (2) an evaluation of different data formats, and (3) visual examples.
We provide pseudocode for generating Sentinel-2 RGB images and NDVI maps from Sentinel-2 L2A in Listing LABEL:lst:rgb and LABEL:lst:ndvi. The RGB algorithm first scales very dark and bright pixels before mapping the image to a uint8 range. We refer to Subsection 3.2 for a detailed description.
@@
import numpy as np def transform_rgb(rbg): Q2, Q98 = np.quantile(rbg, [0.02, 0.98]) rbg = np.where(rbg >= Q2, rbg, Q2 + (rbg - Q2) * 0.5) rbg = np.where(rbg <= Q98, rbg, Q98 + (rbg - Q98) * 0.5) Q02, Q50, Q998 = np.quantile(rbg, [0.002, 0.5, 0.998]) U = max(2000, Q998) L = 0 if Q50 < 1000 else Q02 rbg = (rbg - L) / (U - L) * 255 rbg = np.clip(rbg, 0, 255) return rbg.astype(np.uint8)
@@
def calculate_ndvi(b04, b08): b04 = (b04).clip(0) b08 = (b08).clip(0) ndvi = (b08 - b04) / (b08 + b04 + 1e-6) return ndvi
We compare different dataset formats in terms of data loading speed and storage requirements in Table 4, using the validation split. The experiments were performed on an OpenShift cluster with an average expected I/O speed of 12.5GB/s between data storage and compute nodes. The data loading speed is averaged over 1000 batches using eight CPUs and workers, excluding the initial batches due to slower processing times. The data size is extrapolated based on the size of the validation split.
We evaluate the performance of WebDataset, with individual samples stored as Zarr Zip, Numpy, or TIF files in tar archives. Additionally, we test Zarr Zip files containing batches of 64 samples as utilized by SSL4EO-S12 v1.1 [3]. Zarr Zip files prove to be the most efficient for storage space, while WebDataset with Numpy files offers the fastest I/O speeds with eigth workers (3.6 GB/s). However, Numpy and TIF formats result in substantial storage needs, requiring up to 41.3 TB for the entire dataset. Combining WebDataset with Zarr Zip files optimizes storage space and data loading speed. With 16 workers, WebDataset with Zarr Zip reaches a speed of 0.047 s/batch outperforming Numpy files with 0.054 s/batch as I/O becomes more important than CPU processing. However, further adding CPUs leads to limited improvement. Notably, Hugging Face restricts hosting large-scale datasets to WebDataset or parquet formats. Further experiments with parquet files demonstrated strictly slower data loading times compared to WebDataset.
Table 4: Comparison of data loading speed averaged over 1000 batches with a batch size of 64. WebDataset with Zarr Zip files per sample best optimizes loading speed and storage requirements.
Mod. Format Speed Data size
All WebDataset w/ Zarr 0.096 s/batch 15.4 TB
WebDataset w/ Npy 0.073 s/batch 37.9 TB
WebDataset w/ TIF 0.172 s/batch 41.3 TB
Zarr Zip 0.122 s/batch 14.2 TB
S-2 L2A WebDataset w/ Zarr 0.033 s/batch 5.5 TB
WebDataset w/ Npy 0.022 s/batch 14.0 TB
WebDataset w/ TIF 0.047 s/batch 14.0 TB
Zarr Zip 0.026 s/batch 5.3 TB
S-1 RTC WebDataset w/ Zarr 0.007 s/batch 1.5 TB
WebDataset w/ Npy 0.005 s/batch 2.2 TB
WebDataset w/ TIF 0.019 s/batch 4.2 TB
Zarr Zip 0.009 s/batch 1.4 TB
On the following pages, we present additional examples from both data sources in Figure 6 and 7. All examples are taken from each subset’s first validation tar file, and we manually select representative samples to showcase the full range of landscapes. The focus on urban areas in SSL4EO-S12 is visible, which benefits the pre-training compared to homogeneous sampling. Because of the additional scaling for S-2 RGB, images with few clouds are not very dark, while the clouds keep their structure and are not flattened to a single white value as it is the case with clipping to a max value. Notice the alignment of the cloud and ice classes in the LULC data due to the augmentation using the SEnSeI v2 cloud and ice masks. We only annotate thick clouds as the land surface is still visible below thin cloud layers.
Figure 6: Additional examples from the MajorTOM-Core subset with seven spatiotemporal aligned modalities. Sentinel-2 L2A uses IRRG pseudo-coloring, and Sentinel-1 RTC is visualized in dB scale as VH-VV-VV/VH. Copernicus DEM is scaled based on the image value range with an additional 10 meter buffer.
Figure 7: Additional examples from the SSL4EO-S12 subset with seven spatiotemporal aligned modalities. Sentinel-2 L2A uses IRRG pseudo-coloring, and Sentinel-1 RTC is visualized in dB scale as VH-VV-VV/VH. Copernicus DEM is scaled based on the image value range with an additional 10 meter buffer.